• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Higinbotham

    Log entry time 14:17:13 on April24,2009

    Entry number 267894

    keyword=Beam Current Calibration

    With BCM logger run numbers /usr/opdata/bcm_hallA/BcmLog_14*

    After MCC got everything setup (i.e. got the beam well centered on FC#1 and #2)

    First we did FC#2 in and out at 30 uA.

    BcmLog_146

    Left Arm Run 1338

    Right Arm Run 20218

    -

    Next we put FC#2 and FC#1 in and started BcmLog_147. After 90 sec.,
    we removed FC#1 and took various currents into FC#2.

    -

    Next we again put in both cups and started BcmLog_148. AFter 90 sec., pulled FC#1; after another 90 pulled FC#2 and took beam to the Hall. Then we would put in FC#2 for 90 sec. and lower the current and then send more beam to the hall.

    Runs 1339 and 20219 were taken during this period along with the BcmLog_148.